OpenStudy (anonymous):

Can someone please explain how the solution of the equation below came to be 65536?

OpenStudy (anonymous):

\[e^(8(\ln(4))\]

OpenStudy (anonymous):

We know multiplication is commutative, so: \[ e^{8\ln(4)}=e^{\ln(4)\cdot8}=\left(e^{\ln(4)}\right)^{8}=(4)^8=65536 \]

OpenStudy (anonymous):

(Sadly this isn't the case for complex numbers, but, nevertheless)

OpenStudy (anonymous):

Awwww, thank you so much @LolWolf